# Spatially parallel decoding for multi-qubit lattice surgery

### v2

@@ I Introduction
Here we focus on the surface code [13, 18, 25], a popular QEC that tolerates relatively high physical noise, supports relatively easy logical operations, and only requires nearest-neighbor grid connectivity. In recent years, academic and industry labs have experimentally demonstrated small instances of a surface code logical memory [1, 26, 40]. While an offline decoder — applied after the experiment has completed — is sufficient for such demonstrations, applications with nontrivial information processing will require real time decoding [18].
@@
Whichever decoder one chooses, a single monolithic instance will inevitably struggle to meet the decoding demands that arise during multi-qubit logic gates. For surface codes, lattice surgery [23, 17] is the most efficient way to perform multi-qubit logical operations. When lattice surgery is performed on two or more (possibly distant) patches of logical qubits, they are merged into a single patch for multiple measurement cycles before being split apart; a snapshot of the procedure is shown in Figure 1. It’s worth noting that current state of the art decoders can only meet the throughput demands on a patch that is the size of an individual logical qubit up to distance $20$ to $30$ [2]. Worse, their throughput is inversely proportional to the size of the patch [21]. Meanwhile, lattice surgery requires merging multiple such patches [27] and decoding the merged patch.
@@ II-B Real time decoding with FPGA
For concreteness, we explore the implications of using an FPGA for the inner decoder. FPGAs offer an attractive alternative to general-purpose CPUs for tasks with strict timing requirements and modest programmability requirements. An FPGA is made up of a “fabric” consisting of a great many components whose behaviors are individually programmable, whose input and output can be flexibly routed among other components, and whose timing is uniformly specified so as to support easy latching of inter-component signals. These properties imbue FPGAs with excellent parallelism and pipelinability, enabling them to achieve excellent throughput on high-bandwidth tasks. The primary trade for these abilities is that the FPGA fabric is almost always programmed “once and for all”—that is, the individually programmable components nonetheless have their behaviors fixed for the lifetime of an application.
@@ III Related work
The work in [4] is the only one prior to ours that targets lattice-surgery style fault-tolerant blocks. Compared to it, we put more emphasis on the constraints of hardware accelerators — which are likeliest to meet the throughput requirements. For example, we force the windows to have pre-assigned positions and layers. Additionally, both papers include numerical results on how the buffer width affects decoding accuracy, but we provide more analysis and also study how the influence depends on the physical noise level.
@@ IV-A Connecting decoder modules
Figure 4: Decoding graphs for $X$ errors, used by window A and B respectively. The buffer is marked in pink. In (a), the red edges are the ones that connect the nodes in A’s commit region to the buffer. The two nodes where artificial defects could be introduced are marked by solid black dots. Corrections along the dotted edges are on nodes in the buffer, and are not committed by window A.
@@ IV-C Window configuration for hardware decoding
Figure 6: A window configuration with staggered squares. (a) Only showing the commit region of each window. Each arrow represents a link that communicates information on artificial defects. (b-d) The windows in the first, second, and third layers, respectively. The buffers are shown in lighter color and surrounded by dotted edges. The windows on the last layer do not have buffers. The buffer width in this figure is chosen to improve readability.
@@ V-B Logical errors along the short edge
How long do these strings need to be? For a connected error string (that is correctable by some global decoder) to result in a logical error along the short edge, it should satisfy three conditions:
@@
Let $l$ be the weight of the error string (its number of edges). The right end of the error string can be just above or just below the middle of the patch, and the left end can be just within the commit region of A. The weight of the wrong matching in A, described above, is the sum of the distance from the right end to the rough boundary out of the buffer, and from the left end to the top/bottom boundary. Assuming an odd $d$, the first part is $\max(\frac{d-1}{2}+1-l,0)$, and the second part is $\max(w+1-l,0)$, for a diagonal error string similar to the ones in Figure 10. The wrong matching might be output by decoder A if it has no higher weight than any correct one. Then for $d=15$, a string with $l=4$ ($5,6,7$) can confuse a parallel decoder with $w\leq 3$ ($w\leq 6,w\leq 9,w\leq 12$). Note that when $w\geq d-2$, a connected error string that can confuse the parallel decoder is also long enough to confuse any global decoder.
@@ V-D Generalizing to 2D configurations
As mentioned in Section IV-C, some lattice surgery operations involve thicker merged patches. This section investigates whether the window configuration in Figure 6 works on these patches, and the buffer width requirement for decoding on these patches.
@@
Figure 13: Setups for simulations with 5 windows stacked in 2 rows. The area of the underlying patch is enclosed by the pink square. The vertical and horizontal logical operators are in green and red. In (a) the two original patches that are merged are labelled b1 and o1, while in (b) they are labelled g1 and b1
We imagine a setting where two vertically-stacked square patches, each with the same size as a commit region in Figure 6, are first deformed into rectangles with longer horizontal edges, so that they can perform $Y\otimes Y$ logical measurement [28, 6]. Then the top and bottom rectangles are merged. We perform simulations with the 2 settings shown in Figure 13, where the underlying patch is at the top right and top left corners of the windows in Figure 6, respectively. For baseline we take the LER of the two isolated patches, either before all the logical operations (when they are squares), or right before the merge (when they are longer rectangles).
@@
Figure 14: LER v.s. buffer width at $p=0.4\%$ over 15 rounds, using the setups in Fig. 13a and 13b respectively, with horizontal and vertical logical operators. The size of each smaller square (e.g. the commit region of b1) is chosen to be $15\times 15$. The purple (red) baseline is the LER of two isolated patches, each with size $15\times 29$ ($15\times 15$), taken over 15 rounds and with logical errors along the edge with $d=15$.
The simulation results in Figure 14 shows that, in this setting, the LER of parallel windows is suppressed exponentially with $w$, and is below the baseline before $w$ is grown to more than half the commit region width. Compared to the linear case presented earlier, the parallel windows can be seen to make use of the extra distance present during this logical operation — the buffers extend in both directions rather than one direction. The difference between Figure 14a and b is due to the order that the windows act. In (a), the case with vertical logical operator has higher errors because the intersection of the merged patch and the commit regions of o2 and b3 have short horizontal edges, and o2 and b3 are scheduled before their neighbor g2. This does not apply to (b) because in the lower row of (b), the middle window acts first.
@@ V-E Implications
We conclude that a sufficient buffer width $w$ is needed for achieving high accuracy. Since linear settings (Figure 7) have larger buffer width requirements than multi-row settings (Figure 13), it suffices to choose $w$ based on Section V-B. Section V-C shows that the size of the windows (or specifically, the intersection of a patch with a window) also bound the fidelity that can be achieved. This is acceptable as long as a window is wider than an isolated patch of code. One mapping/compiling constraint should be noted though. When a patch spans multiple commit regions, and its intersection with one of the windows is too small, the type of logical errors in Figure 12 can be large. This is the case in Figure 13a, where the intersection of the patch with the commit region of o2 has width $d/2$. When $w$ is at least half of $d$, o2’s total width is at least $d$, so the contribution from this type of error is not too large. However, the intersection of a patch with a commit region should not be smaller than half the commit region.

### v3

@@ I Introduction
Here we focus on the surface code [13, 19, 26], a popular QEC that tolerates relatively high physical noise, supports relatively easy logical operations, and only requires nearest-neighbor grid connectivity. In recent years, academic and industry labs have experimentally demonstrated small instances of a surface code logical memory [1, 27, 43] and lattice surgery [16]. While an offline decoder — applied after the experiment has completed — is sufficient for such demonstrations, applications with nontrivial information processing will require real time decoding [19].
@@
Whichever decoder one chooses, a single monolithic instance will inevitably struggle to meet the decoding demands that arise during multi-qubit logic gates. For surface codes, lattice surgery [24, 18, 28, 29, 30, 36] is the most efficient way to perform multi-qubit logical operations. When lattice surgery is performed on two or more (possibly distant) patches of logical qubits, they are merged into a single patch for multiple measurement cycles before being split apart; a snapshot of the procedure is shown in Figure 1. It’s worth noting that current state of the art decoders can only meet the throughput demands on a patch that is the size of an individual logical qubit up to distance $20$ to $30$ [2]. Worse, their throughput is inversely proportional to the size of the patch [22]. Meanwhile, lattice surgery requires merging multiple such patches [29] and decoding the merged patch.
@@ II-B Real time decoding with FPGA
For concreteness, we explore the implications of using an FPGA for the inner decoder. FPGAs offer an attractive alternative to general-purpose CPUs for tasks with strict timing requirements and modest programmability requirements. An FPGA is made up of a “fabric” consisting of a great many components whose behaviors are individually programmable, whose input and output can be flexibly routed among other components, and whose timing is uniformly specified so as to support easy latching of inter-component signals. These properties imbue FPGAs with excellent parallelism and pipelinability, enabling them to achieve excellent throughput on high-bandwidth tasks. The primary trade-off for these abilities is that the FPGA fabric is almost always programmed “once and for all”—that is, the individually programmable components nonetheless have their behaviors fixed for the lifetime of an application.
@@ III Related work
A paper from PsiQuantum [4] is the only prior work that targets lattice-surgery style fault-tolerant blocks. Its focus is different from our work though. It prioritizes minimizing the latency of a software implementation and does not address “further hardware and systems considerations that are relevant”. However, we put more emphasis on the constraints of hardware accelerators, and consider throughput instead of latency as the more important measure of speed for real time decoders. For example, they develop algorithms that take lattice surgery blocks as input, then assign windows, while we explicitly require that windows have pre-assigned positions so that they are compatible with hardware accelerators. Both papers include numerical results on how the buffer width affects decoding accuracy, but we provide more analysis and also study how the influence depends on the physical noise level.
@@ IV-A Connecting decoder modules
Figure 4: Decoding graphs for $X$ errors, used by (a) window A and (b) window B. In both (a) and (b), the buffer is marked in pink, and the two nodes where artificial defects could be introduced are marked by solid black dots. In (a), the red edges are the ones that connect the nodes in A’s commit region to the buffer. Corrections along the dotted edges are on nodes in the buffer, and are not committed by window A. In (b) the decoding graph has solid edges in the buffer, because the corrections in the buffer are finalized in this step.
@@ IV-C Window configuration for hardware decoding
Figure 6: A window configuration with staggered squares. (a) Only showing the commit region of each window. Each arrow represents a one-way link that communicates information on artificial defects. (b-d) The windows in the first, second, and third layers, respectively. The buffers are shown in lighter color and surrounded by dotted edges. The windows on the last layer do not have buffers. The buffer width in this figure is chosen to improve readability.
@@ V-B Logical errors along the short edge
How long do these strings need to be? For a connected error string (that is correctable by some global decoder) to result in a logical error along the short edge, it should satisfy all the three conditions below:
@@
Let $l$ be the weight of the error string (its number of edges). The right end of the error string can be just above or just below the middle of the patch, and the left end can be just within the commit region of A. The weight of the wrong matching in A, described above, is the sum of the distance from the right end to the rough boundary out of the buffer, and from the left end to the top/bottom boundary. Assuming an odd $d$, the first part is $\max(w+1-l,0)$, and the second part is $\max(\frac{d-1}{2}+1-l,0)$, for a diagonal error string similar to the ones in Figure 10. The wrong matching might be output by decoder A if it has no higher weight than any correct one. Then for $d=15$, a string with $l=4$ ($5,6,7$) can confuse a parallel decoder with $w\leq 3$ ($w\leq 6,w\leq 9,w\leq 12$). Note that when $w\geq d-2$, a connected error string that can confuse the parallel decoder is also long enough to confuse any global decoder.
@@ V-D Generalizing to 2D configurations
As mentioned in Section IV-C, some lattice surgery operations involve wider merged patches. This section investigates whether the window configuration in Figure 6 works on these patches, and the buffer width requirement for decoding on these patches.
@@
Figure 13: Setups for simulations with 5 windows stacked in 2 rows. As in Figure 6, the color of a window indicates the step that the decoder acts on it. The area of the underlying patch is enclosed by the pink square. The vertical and horizontal logical operators are in green and red. In (a) the two original patches that are merged are labelled b1 and o1, while in (b) they are labelled g1 and b1.
We perform simulations with the 2 settings shown in Figure 13, where we allocate an underlying patch in different regions of Figure 6, and perform decoding with the given window configurations. In practice, such a large patch could be the result of a merge: we start with two vertically-stacked square patches, each with the same size as a commit region in Figure 6. They are first deformed into rectangles with longer horizontal edges, so that they can perform $Y\otimes Y$ logical measurement [30, 6]. Then the top and bottom rectangles are merged. This example scenario motivates our choices of baselines: the total LERs of the two isolated patches, either before all the logical operations (when they are squares), or right before the merge (when they are longer rectangles).
@@
Figure 14: LER v.s. buffer width at $p=0.4\%$ over 15 rounds, using the setups (a) in Fig. 13a, and (b) in Fig. 13b. We show the different logical errors from using the horizontal and the vertical logical operators. The size of each smaller square (e.g. the commit region of b1) is chosen to be $15\times 15$. The blue (purple) baseline is the LER of two isolated patches, each with size $15\times 29$ ($15\times 15$), taken over 15 rounds and with logical errors along the edge with $d=15$.
The simulation results in Figure 14 shows that, in this setting, the LER of parallel windows is suppressed exponentially with $w$, and is below the baseline before $w$ is grown to more than half the commit region width. Compared to the linear case presented earlier, the parallel windows can be seen to make use of the extra distance present during this logical operation — the buffers extend in both directions rather than one direction. The difference between Figures 14a and b is due to the order in which the windows act. In (a), the case with vertical logical operator has higher errors because the intersection of the merged patch and the commit regions of o2 and b3 have short horizontal edges, and o2 and b3 are scheduled before their neighbor g2. This does not apply to (b) because in the lower row of (b), the middle window acts first.
@@ V-E Implications
The results in this section could help us determine the buffer width $w$ that is needed for achieving high accuracy. Linear settings (Figure 7) have larger buffer width requirements than multi-row settings (Figure 13), and the logical error rate in the linear setting is dominated by errors along the shorter edge. Therefore, it suffices to choose $w$ based on one’s error budget and the physical noise level, according to the results in Section V-B. In general, $w$ should be around $d/2$. Its size could be larger if the physical noise level is low or if a closer match to the logical error rate of the global decoder is required. Conversely, it could be smaller if the physical noise level is high or if a looser match is acceptable.
Section V-C shows that the size of the windows (or specifically, the intersection of a patch with a window) also bound the fidelity that can be achieved. This is acceptable as long as a window is wider than an isolated patch of code. One mapping/compiling constraint should be noted though. When a patch spans multiple commit regions, and its intersection with one of the windows is too small, the type of logical errors in Figure 12 can be large. This is the case in Figure 13a, where the intersection of the patch with the commit region of o2 has width $d/2$. When $w$ is at least half of $d$, o2’s total width is at least $d$, so the contribution from this type of error is not too large. However, the intersection of a patch with a commit region should not be smaller than half the commit region.
@@ VII Conclusion
Acknowledgements
This work is funded in part by the STAQ project under award NSF Phy-1818914/232580 and in part by the NSF Quantum Leap Challenge Institute for Hybrid Quantum Architectures and Networks (NSF Award 2016136).